# Proximal Interacting Particle Langevin Algorithms

### v2

@@ 4.5 Algorithm comparison
Theorem 4.1 and 4.2 and Theorem B.1 in Appendix B permit the following bounds for $\mathbb{E}\left[\|\theta_{n}^{N}-\bar{\theta}^{\star}\|^{2}\right]^{1/2}=\mathcal{O}(\varepsilon)$ in terms of the key parameters $d_{\theta},d_{x}$
$\lambda$ $N$ $\gamma$ $n$
where $\delta>0$ is any small positive constant.
We point out that even if all algorithms allow for the same bound w.r.t. $\gamma$, the bound for MYPGD in Theorem B.1 requires more stringent assumptions on $\gamma$: while Theorem 4.1 requires $\gamma_{0}<\min\{(L_{g_{1}}+\lambda^{-1})^{-1},2\mu^{-1}\}$, Theorem B.1 requires $\gamma_{0}<(L_{g_{1}}+\lambda^{-1}+\mu)^{-1}$ which is smaller than $\min\{(L_{g_{1}}+\lambda^{-1})^{-1},2\mu^{-1}\}$. PIPGLA allows for a less restrictive choice of $\gamma<L_{g_{1}}$, but requires more stringent assumptions on $\lambda$.
@@
We can compare the algorithms in terms of their computational requirements too. We account for the computational cost of running each algorithm for $n$ iterations with $N$ particles and time discretisation step $\gamma$, while guaranteeing an $\mathcal{O}(\varepsilon)$ error, in terms of component-wise evaluations of $\nabla g_{1}$ and $\operatorname{prox}_{g_{2}}^{\lambda}$, and sampling of independent standard $1$-dimensional Gaussians.
5 Numerical Experiments
We evaluate the numerical performance of the proximal interacting particle algorithms proposed in Section 3 by applying them to train a Bayesian logistic regression model, a Bayesian neural network for MNIST classification and a sparse matrix completion model.
5.1 A hierarchical model
As a first example, we consider a hierarchical model involving a single scalar unknown parameter $\theta$, $d_{x}$ i.i.d. latent random variables following a distribution inspired by the elastic net. The prior $p_{\theta}(x)$ is defined as a product of the Gaussian and the Laplace distribution, with mean parameter $\theta$ [Li and Lin, 2010], that is,
$p_{\theta}(x_{i})\propto e^{-\beta_{1}|x_{i}-\theta|-\beta_{2}(x_{i}-\theta)^{2}},\;\beta_{1},\beta_{2}\geq 0.$
We consider a Gaussian likelihood $p(y_{i}|x_{i})=\mathcal{N}(y_{i};x_{i},1)$ which results in the joint log-likelihood:
$-\log p_{\theta}(x,y)$ $=\underbrace{\beta_{1}\sum_{i=1}^{d_{x}}|x_{i}-\theta|}_{g_{2}(\theta,x)}+\underbrace{C+\beta_{2}\sum_{i=1}^{d_{x}}(x_{i}-\theta)^{2}+\sum_{i=1}^{d_{x}}\frac{(y_{i}-x_{i})^{2}}{2}}_{g_{1}(\theta,x)},$
where we identify the functions $g_{1}$ and $g_{2}$ as above. In the standard posterior sampling setting, potentials similar to the one considered here appear in Bayesian inverse problems [Pereyra, 2016; Durmus et al., 2018; Crucinio et al., 2023].
Note that the function $g_{1}$ is gradient Lipschitz and lower bounded by $C$, and $g_{2}$ is proper, convex, lower semi-continuous, lower bounded and Lipschitz. If we take $\beta_{2}$ bounded away from zero, i.e., $\beta_{2}\geq\varepsilon>0$, then $g_{1}$ is $\mu$-strongly convex, where $\mu$ is given by be the lowest eigenvalue of the Hessian matrix of $g_{1}$ and depends on $\beta_{2}$. Besides, in the set where $g_{2}$ is differentiable it satisfies $\|\nabla_{\theta}g_{2}\|=\|\nabla_{x}g_{2}\|=1$ and ($\|\nabla^{2}_{(\theta,x)}g_{2}\|=0$. Therefore, assumption A5 is also satisfied.
We derive the proximal mapping for $g_{2}$ in Appendix D.1.1. Table 2 shows the normalised MSEs (NMSE) for $\theta$, $\|\theta_{K+1}^{N}-\theta\|^{2}/\|{\theta}\|^{2}$. Each algorithm is run 100 times using different starting points for 50 and 500 particles and 2500 steps. The NMSEs and computation times provided are averaged over the 100 replicates. The values of $\lambda$ and $\gamma$ are the optimal values found by performing a grid search. The execution times for the three algorithms MYIPLA, PIPGLA and MYPGD are very similar. The performance of MYIPLA and MYPGD is better than that of PIPGLA for $N=50$ particles while PIPGLA clearly outperforms the others for $N=500$ particles. Note also how the NMSE decreases when increasing the number of particles for PIPGLA and MYPGD.
Table 1: Toy hierarchical example. Normalised MSE (NMSE) for $\theta$ for different algorithm when run 100 times using 50 and 500 particles, 2500 steps and different starting points. Computation times and NMSEs are averaged over the 100 replicates.
Algorithm NMSE (×$\times 10^{-4}$) Times (s) $\lambda$ $\gamma$
$N=50$ $N=500$ $N=50$ $N=500$
MYIPLA $\mathbf{1.98\pm 3.27}$ $3.06\pm 0.74$ $31\pm 12$ $\mathbf{170\pm 23}$ $0.25$ $0.03$
PIPGLA $10.74\pm 13.35$ $\mathbf{0.25\pm 0.12}$ $36\pm 15$ $185\pm 31$ $0.005$ $0.05$
MYPGD $3.62\pm 4.81$ $3.59\pm 1.60$ $\mathbf{27\pm 10}$ $175\pm 22$ $0.25$ $0.05$
5.2 Bayesian logistic regression
Here we analyse a Bayesian logistic regression latent variable model. We consider a similar set-up to the one described in De Bortoli et al. [2021] and employ a synthetic dataset consisting of $d_{y}=900$ datapoints (see Appendix D.2 for details). The latent variables are the $d_{x}=50$ regression weights, to which we assign an isotropic Laplace prior $p_{\theta}(x)=\prod_{i=1}^{d_{x}}\text{Laplace}(x_{i}|\theta,1)$ or a uniform prior $p_{\theta}(x)=\prod_{i=1}^{d_{x}}\mathcal{U}(x_{i}|-\theta,\theta)$. The likelihood is given by $p_{\theta}(y|x)=\prod_{j=1}^{d_{y}}s(v_{j}^{T}x)^{y_{j}}s(-v_{j}^{T}x)^{1-y_{j}}$, where $v_{j}\sim\mathcal{U}(-1,1)^{\otimes d_{x}}$ are a set of synthetic $d_{x}$-dimensional covariates sampled from a uniform distribution and $s(u):=e^{u}/(1+e^{u})$ is the logistic function. The true value of $\theta$ is set to $\theta=-4$ for the Laplace prior and $\theta=1.5$ for the uniform one.
and for the uniform prior, we obtain
where we introduce $g_{1}$ and $g_{2}$ as above and $\imath_{\mathcal{K}}$ is the convex indicator of $\mathcal{K}$ defined by $\imath_{\mathcal{K}}(x)=0$ if $x\in\mathcal{K}$ and $\imath_{\mathcal{K}}(x)=\infty$ otherwise.
In both cases we have that
where $C$ is a constant. As shown in Akyildiz et al. [2023, Section 6.1.1], the function $g_{1}$ is gradient Lipschitz and strictly convex but not strongly convex. The function $g_{2}$ satisfies A1 for both the Laplace and the uniform prior, as observed in Pereyra [2016], in the case of the Laplace prior $g_{2}$ also satisfies A3 while the uniform prior does not lead to a Lipschitz $g_{2}$. Since $g_{1}$ does not depend on $\theta$, A5 holds for the Laplace prior as shown in the previous example.
5.2.1 Comparing approximations of $\operatorname{prox}_{g_{2}}^{\lambda}$
We derive approximations of the proximal mapping for $g_{2}$ for both the Laplace prior and the uniform prior in Appendices D.1.1 and D.1.3, respectively. The exact solution for the proximal operator for these two priors is not available in closed form, however Schmidt et al. [2011]; Ehrhardt et al. [2023] provide convergence guarantees when the minimisation in Definition 2.1 can only be solved up to a certain accuracy. In particular, they show that the error in the proximity operator calculation is controlled in an appropriate way and, in some cases, inexact proximal-gradient strategies can attain the same convergence rates as exact methods. The proximity map for PIPULA and PPGD (i.e. when $\lambda=\gamma$) is derived in D.1.4
Figure 1: Convergence rate of the variance of the parameter estimates produced by MYIPLA and PIPGLA over 100 Monte Carlo runs for different initialisations each $N\in{10,100,1000}$. We can see that the $\mathcal{O}(1/N)$ convergence rate holds for the second moments as suggested by our results in Section 4.
We consider both an iterative approach [Parikh and Boyd, 2014, Section 2.3] and an approximation of $\operatorname{prox}_{g_{2}}^{\lambda}$. Since the true value of $\theta$ is known, we use this experiment to validate the proposed algorithms (and their variations) and to investigate the effect of an approximate proximity map against one obtained by an iterative fixed-point procedure for the Laplace prior. In the case of the uniform prior, we only consider an approximation to the proximity map.
Figure 1 shows the variance of the $\theta$ estimates produced by MYIPLA and PIPGLA computed over 100 Monte Carlo runs with different initialisations $\theta_{0}$. We observe that the second moment, $\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}^{\star}\|^{2}]$, decreases with rate $\mathcal{O}(1/N)$ as suggested by Theorems 4.1 and 4.2, and that the iterative algorithms have a slightly lower variance compared to their approximate versions, with PIPGLA having better performance than MYIPLA. It is also important to highlight that for all algorithms considered, approximate solvers are on average 25% faster than iterative solvers (see Table 2).
5.2.2 Overall performances
We compare the performances of MYIPLA, PIPGLA, MYPGD, PIPULA and PPGD through the normalised MSEs (NMSE) for $\theta$, $\|\theta_{K+1}^{N}-\theta\|^{2}/\|{\theta}\|^{2}$. Each algorithm is run 500 times using different starting points for 50 particles and 5000 steps. When using an iterative solver for the proximity map we use 40 iterations at each step. Table 2 provides the NMSEs and computation times which are averaged over the 500 replicates. The values for $\lambda,\gamma$ in the different algorithms are chosen optimally after performing a grid search. The specific values are provided in Appendix D.2, Table 7. In the case of PIPGLA the optimal values for $\lambda,\gamma$ are obtained when $\lambda=\gamma$ as suggested by Salim et al. [2019].
For the Laplace prior, PIPGLA attains the best performance followed by MYIPLA, with the iterative solver being slightly better but considerably slower and therefore computationally more expensive. Therefore, considering a trade-off between accuracy and computational cost, PIPGLA and MYIPLA implemented with the iterative solver outperform the other algorithms.
For the uniform prior, MYPGD demonstrates significantly better performance than all the other algorithms; this is likely due to the lack of diffusive term in the corresponding SDE which is beneficial when dealing with a compactly supported prior. Figure 4 in Appendix D.2 shows that the $\theta$-iterates of MYPGD, MYIPLA, and PIPGLA converge to similar stationary points for the different initial values $\theta_{0}$ although for the last two algorithms the stationary distribution does not coincide with the true parameter value.
@@ 5.2.2 Overall performances
PPGD Approx $14.70\pm 4.42$ $3.63\pm 4.93$ $102.6\pm 5.1$ $107.9\pm 5.5$
Iterative $19.04\pm 1.34$- $122.3\pm 5.1$-
PIPULA Approx $12.18\pm 1.62$ $4.71\pm 6.02$ $98.8\pm 5.7$ $101.0\pm 4.0$
Iterative $19.22\pm 1.28$- $126.2\pm 3.8$-
@@
5.3 Bayesian neural network with sparse prior
To test our algorithms on an example with more complex posteriors, we turn to Bayesian neural networks whose posteriors are notoriously multimodal, where we consider a sparsity-inducing prior on the weights, which is non-smooth and therefore it cannot be handled using the framework of Kuntz et al. [2023]; Akyildiz et al. [2023].
We consider an analogous setting to Yao et al. [2022] and Kuntz et al. [2023], and apply a Bayesian two-layer neural network to classify MNIST images. Similarly to Yao et al. [2022], we avoid the cost of computing the gradients on a big dataset by subsampling 1000 data points with labels 4 and 9. The input layer of the network has 40 nodes and 784 inputs (since we are considering $28\times 28$ images), and the output layer has 2 nodes. The latent variables are the weights, $w\in\mathbb{R}^{d_{w}:=40\times 784}$, of the input layer and those, $v\in\mathbb{R}^{d_{v}:=2\times 40}$, of the output layer. We use tanh activation functions, a softmax output layer, and also simplify the problem by setting all network biases to zero. This scenario is equivalent to assuming that the datapoints’ labels $l$ are conditionally independent given the features $f$ and network weights $x=(w,v)$, and therefore have the following probability density
$p_{\theta}(x,\mathcal{Y}_{\text{train}})=\prod_{i}\text{Laplace}(w_{i}|0,e^{2\alpha})\prod_{j}\text{Laplace}(v_{j}|0,e^{2\beta})\prod_{(f,l)\in\mathcal{Y}_{\text{train}}}p(l|f,x).$
We note that the log density can be decomposed as
$-\log p_{\theta}(x,\mathcal{Y}_{\text{train}})=\underbrace{2d_{w}\alpha+\sum_{i}|w_{i}|e^{-2\alpha}+2d_{v}\beta+\sum_{j}|v_{j}|e^{-2\beta}}_{g_{2}(\theta,x)}\underbrace{-\sum_{(f,l)\in\mathcal{Y}_{\text{train}}}\log p(l|f,x)}_{g_{1}(\theta,x)},$
where $d_{w}$ and $d_{v}$ denote the dimensions of the weights $w$ and $v$, respectively, $g_{1}$ is differentiable and does not depend on $\theta$ and $g_{2}$ is proper, convex and lower semi-continuous. We derive an approximation to the proximity map of $g_{2}$ in Appendix D.1.2.
To allow a fair comparison with other algorithms, we use the same performance metrics as in Kuntz et al. [2023]; see Appendix D.3 for their precise definition. We consider the average classification error over the test set $\mathcal{Y}_{\text{test}}$, i.e.
$\text{Error}:=\frac{1}{|\mathcal{Y}_{\text{test}}|}\sum_{(f,l)\in\mathcal{Y}_{\text{test}}}\mathds{1}\{l=\hat{l}(f)\},$
where $\hat{l}(f)$ denotes the most likely label for feature $f$ (i.e. the $\operatorname*{arg\,max}$ of the approximation of the posterior predictive distribution at $\bar{\theta}_{\star}$), and the log pointwise predictive density (LPD, Vehtari et al. [2017])
$\textrm{LPD}:=\frac{1}{|\mathcal{Y}_{\text{test}}|}\sum_{(f,l)\in\mathcal{Y}_{\text{test}}}\log\left(\frac{1}{N}\sum_{i=1}^{N}p(l|f,X_{K+1}^{i})\right),$
where $X_{K+1}^{i}$ are the samples of the final particle cloud.
(a) Laplace prior. Algorithm MYIPLA
(b) Normal prior. Algorithm IPLA
(c) Laplace prior. Algorithm IPLA
(d) Laplace prior. Algorithm PIPGLA
Figure 2: Histogram and density estimation of the weights of a BNN for a randomly chosen particle from the final (500 steps) cloud of 100 particles.
5.3.1 Laplace vs Normal prior
One may ask whether the Laplace prior is more appropriate in this setting than the Normal one. Jaynes [1968] provides two reasons why the Laplace prior is particularly suitable for Bayesian neural network models. Firstly, for any feedforward network there is a functionally equivalent network in which the weight of a non-direct connection has the same size but opposite sign, therefore consistency demands that the prior for a given weight $w$ is a function of $|w|$ alone. Secondly, if it is assumed that all that is known about $|w|$ is its scale, and that the scale of a positive quantity is determined by its mean rather than some higher order moment, then the maximum entropy distribution for a positive quantity constrained to a given mean is the exponential distribution. It would follow that the each signed weight $w$ has a Laplace density [Williams, 1995].
We analyse if the sparsity-inducing nature of the Laplace prior has in practice this effect on the value of the final weights in the case of MYIPLA. For this, we set $N=100$ and observe the distribution of the weights for a randomly chosen particle from the final particle cloud $X_{500}^{1},\dots,X_{500}^{N}$ and compare it to that obtained with a Normal prior (Figure 2). We note that our experiment (Fig. 2(a)) leads to final weights with values highly concentrated around zero in comparison to the normal prior (Fig. 2(b)) where the values of the weights are more spread out, leading to a posterior distribution with heavier tails (see also the standard deviation column in Table 4).
The sparse representation of our experiment also has the advantage of producing models that are smaller in terms of memory usage when small weights are zeroed out. To investigate this, we set to zero all weights below a certain threshold and analyse the performance of the compressed weight matrices. We consider two cases, averaging the particles of the final cloud $X_{500}^{1},\dots,X_{500}^{100}$, applying the threshold and then calculating the performance, and secondly, setting to zero small values of each particle of the cloud and averaging the performance of each particle. We compare the results for the Bayesian neural networks with Laplace and Normal priors (Table 3). It is important to note that when applying the same threshold to both cases, the Laplace prior leads to a very compressed weight matrix compared to the Normal prior, i.e. there is a significant difference in the percentage of weights set to zero. We observe that when setting the same proportion of weights to zero in both layers, the performance of the BNN with Laplace priors is better in terms of the log pointwise predictive density than that of the BNN with Normal priors, especially when averaging the final cloud of particles before computing the performance. Further results for other proposed algorithms are presented in Appendix D.3.
@@ 5.3.1 Laplace vs Normal prior
particles? Layer 1 Layer 2 Layer 1 Layer 2
Laplace $74$ $48$ $0.2$ $0.2$ $\mathbf{7.0}$ $\mathbf{-0.23}$
$56$ $35$ $1$ $1$ $\mathbf{1.5}$ $\mathbf{-0.07}$
Normal $74$ $48$ $0.5$ $1.1$ $15$ $-0.74$
$16$ $15$ $0.2$ $0.2$ $16$ $-0.78$
$56$ $35$ $7$ $4$ $2.0$ $-0.11$
$8.6$ $7.1$ $1$ $1$ $1.5$ $-0.10$
5.3.2 Comparing proximal algorithms with IPLA
We compare the performance of MYIPLA, MYPGD and PIPGLA with IPLA [Akyildiz et al., 2023] obtained by ignoring the non-differentiability of the model density. Adapting the results of Johnston and Sabanis [2023], one can show that IPLA can be analysed within this setting and would have a similar convergence guarantee to that given in Akyildiz et al. [2023]. In this section, we show that, even with a Laplace prior, the IPLA implementation does not produce sufficient sparsity, compared to the methods we propose in this paper.
These results are summarised in Table 4 when using 50 particles, 250 steps and choosing $\lambda$ and $\gamma$ optimally after performing a grid search. The table also includes computation times and the standard deviation of the weight matrix $w$ for a randomly chosen particle in the final cloud. This latter value is included to asses whether the model is able to induce sparsity in the computed weights. We observe that while the performance of all the algorithms is similar, the standard deviation of the weights for the non-proximal method, IPLA, is significantly higher compared to the proximal algorithms. This results in posterior distributions for the weights which are less concentrated (see, e.g., Figure 2(c)) and thus the sparsity-inducing effect of the Laplace prior is reduced by the use of algorithms which do not take into account the non-differentiability of the target. We can visually verify that proximal algorithms lead to concentrated posterior distributions, see, e.g., Figures 2(a) and 2(d). In particular when using PIPGLA, 70% and 12% of the weights of $w$ and $v$ of the final particle cloud are exactly zero up to machine precision.
Table 4: Bayesian neural network. Test errors and log pointwise predictive density (LPD) achieved using the final particle cloud $X_{250}^{1},\dots,X_{250}^{N}$ for different algorithm when using 50 particles. Computation times and standard deviation of the empirical distribution of the weight matrix $w$ for a randomly chosen particle are also provided.

### v3

@@ 4.5 Algorithm comparison
Theorems 4.1,  4.2 and  B.1 (Appendix) allow us to derive complexity estimates for $\lambda$, the number of particles $N$, $\gamma$, and the number of steps $n$ to achieve an error $\mathbb{E}\left[\|\theta_{n}^{N}-\bar{\theta}^{\star}\|^{2}\right]^{1/2}=\mathcal{O}(\varepsilon)$, see blue columns of Table 1. These bounds are expressed in terms of the key parameters $d_{\theta},d_{x}$. Details for deriving these bounds are provided in Appendix D. It is important to mention that although all algorithms yield the same complexity estimates in terms of $\gamma$, MYPGD requires more stringent assumptions on $\gamma$. Specifically, while MYIPLA (Theorem 4.1) requires $\gamma_{0}<\min\{(L_{g_{1}}+\lambda^{-1})^{-1},2\mu^{-1}\}$, MYPDG (Theorem B.1) requires $\gamma_{0}<(L_{g_{1}}+\lambda^{-1}+\mu)^{-1}$, which is strictly smaller. In contrast, PIPGLA allows for a more flexible choice of $\gamma\leq 1/L_{g_{1}}$, but requires stronger assumptions on $\lambda$.
@@
We also compare the algorithms in terms of their computational requirements. In particular, we evaluate the computational cost of running each algorithm for $n$ iterations with $N$ particles and a time discretisation step $\gamma$, while guaranteeing an $\mathcal{O}(\varepsilon)$ error. The comparison is based on the number of component-wise evaluations of $\nabla g_{1}$ and $\operatorname{prox}_{g_{2}}^{\lambda}$, and independent standard $1$-$d$ Gaussians samples. These costs are summarised in the green columns of Table 1.
[…]
Table 1: Comparison between algorithms (Section 4.5). $\delta>0$ is any small positive constant.
[…]
Table 2: Performance of Bayesian logistic regression for Laplace and uniform priors.
@@
5 Experiments
The code is available in https://github.com/paulaoak/proximal-ipla. Additional experiments for low-rank matrix completion are provided in Appendix E.5, where the goal is to recover a low-rank matrix from partially observed and noisy data.
5.1 Bayesian Logistic Regression
We consider a similar set-up to De Bortoli et al. [2021] and employ a synthetic dataset consisting of $d_{y}=900$ datapoints (see Appendix E.2 for details). The latent variables are the $d_{x}=50$ regression weights, to which we assign an isotropic Laplace prior $p_{\theta}(x)=\prod_{i=1}^{d_{x}}\text{Laplace}(x_{i}|\theta,1)$ or a uniform prior $p_{\theta}(x)=\prod_{i=1}^{d_{x}}\mathcal{U}(x_{i}|-\theta,\theta)$. The likelihood is given by $p_{\theta}(y|x)=\prod_{j=1}^{d_{y}}s(v_{j}^{T}x)^{y_{j}}s(-v_{j}^{T}x)^{1-y_{j}}$, where $v_{j}\sim\mathcal{U}(-1,1)^{\otimes d_{x}}$ are a set of $d_{x}$-dimensional covariates sampled from a uniform distribution and $s(u)$ is the logistic function. The true value of $\theta$ is set randomly to $\theta=-4$ for the Laplace prior and $\theta=1.5$ for the uniform one.
Approximations of the proximal mapping for both priors are derived in Appendices E.1.1 and E.1.3, as closed-form solutions are unavailable. We compare these approximations to an iterative approach for computing the proximal map, which is feasible only for the Laplace prior due to instabilities in the uniform case. Figure 1 shows the variance of the $\theta$ estimates produced by MYIPLA and PIPGLA computed over 100 runs for different initialisations (left) and the sequence of $\theta$ estimates for 50 particles (right) in the case of a Laplace prior. We observe that the variance of the parameter estimates decreases with rate $\mathcal{O}(1/N)$ as suggested by Theorems 4.1 and 4.2, and that the iterative algorithms have a slightly lower variance compared to their approximate versions, with PIPGLA having better performance than MYIPLA. It is also important to highlight that for all algorithms considered, approximate solvers are on average 25% faster than iterative solvers (see Table 2). Besides, in Table 2 (expanded in Table 6), we compare the performances of the different proximal algorithms through the normalised MSEs (NMSE) for $\theta$. In the uniform case, MYPGD outperforms all other algorithms; this is likely due to the lack of diffusive term in the corresponding SDE which is beneficial when dealing with a compactly supported prior. While PIPGLA also produces estimates within the support, they exhibit a larger bias.
Figure 1: Laplace prior. Left: convergence rate of the variance of the parameter estimates against $N$ produced by MYIPLA and PIPGLA over 100 runs. We see that the $\mathcal{O}(1/N)$ convergence rate holds for the second moments. Right: evolution of the normalised MSE for 50 particles over 100 runs.
5.2 Bayesian Neural Network
To evaluate our algorithms on complex multimodal posteriors, we consider a Bayesian neural network with a sparsity-inducing prior on the weights for MNIST image classification. Following Yao et al. [2022] and Kuntz et al. [2023], we use a two-layer network with tanh activation functions and avoid the cost of computing the gradients on a big dataset by subsampling 1000 data points with labels 4 and 9. The input layer of the network has 40 nodes and 784 inputs and the output layer has 2 nodes. The latent variables are the weights, $w\in\mathbb{R}^{d_{w}:=40\times 784}$ and $v\in\mathbb{R}^{d_{v}:=2\times 40}$ of the input and output layers, respectively. We assign priors $p_{\alpha}(w)=\prod_{i}\text{Laplace}(w_{i}|0,e^{2\alpha})$ and $p_{\beta}(v)=\prod_{i}\text{Laplace}(v_{i}|0,e^{2\beta})$ and learn $\theta=(\alpha,\beta)$ from the data. One may ask whether the Laplace prior is more appropriate in this setting than the Normal one. Jaynes [1968] shows that the Laplace prior naturally arises for Bayesian neural network models (see Appendix E.3.1 for details). We analyse the sparsity-inducing effect of the Laplace prior by examining the distribution of the weights for a randomly chosen particle from the final particle cloud and comparing it to that obtained with a Normal prior. We note that our methods (Fig. 2(a), 2(b)) lead to final weights with values highly concentrated around zero in comparison to the Normal prior (Fig. 2(c)). The sparse representation of our algorithm has the advantage of producing models that are smaller in terms of memory usage when small weights are zeroed out. This is investigated in Table 8 in the Appendix. Furthermore, we compare the performance of the PIPLA family against IPLA which ignores the non-differentiability of the model density. Figure 2(d) shows that, despite using a Laplace prior, IPLA fails to induce sufficient sparsity compared to our proposed methods. Quantitative results for the variance of the weights and error metrics are shown in Table 1, comparing our approach with other algorithms in the literature. Appendix E.3.2 provides additional results on more complex datasets. In particular, we apply our methods to a classification task using CIFAR10 dataset. Furthermore, in Appendix E.3.3, we also explore the application of our methods to neural networks with non-differentiable activation functions, such as ReLU.
(a) Laplace prior, MYIPLA
(b) Laplace prior, PIPGLA
(c) Normal prior, IPLA
(d) Laplace prior, IPLA
Figure 2: Histogram (blue) and density estimation (red) of the BNN weights for a randomly chosen particle. Our methods (top) produce sparser weights, which is crucial for compressibility, compared to IPLA (bottom), which ignores the non-differentiabilities.
@@ 5.2 Bayesian Neural Network
Table 3: Bayesian neural network. Test errors and log pointwise predictive density (LPPD) achieved using the final particle cloud with $N=50$. Computation times and standard deviation of the empirical distribution of the weight matrix $w$ are also provided.